OpenStudy (anonymous):

Use trigonometric identities to transform the left side of the equation into the right side.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\sin ^2 \theta - \cos ^2\theta = 2 \sin ^2 \theta - 1\]

OpenStudy (anonymous):

i've got it started and i have \[\sin ^2\theta - (1- \sin ^2 \theta) = 2\sin ^2 \theta -1\] idk what to do next...please help!!

OpenStudy (anonymous):

@amy101 thanks for what?

OpenStudy (luigi0210):

Easy, distribute that negative :)

OpenStudy (anonymous):

oh i see!! :) Thank you!

OpenStudy (luigi0210):

You're welcome~